OpenStudy (anonymous):

Multiply: 5xy(2x - y) A. 10xy - 5xy B. 10x^2y - 5xy^2 C. 10x^2 - 5xy + 25y^2 D. 10x^2 - xy + y^2

sam (.sam.):

distribute 5xy(2x - y) \[5xy \times 2x - 5xy \times y=?\]

OpenStudy (anonymous):

is B the answer?

sam (.sam.):

correct :)